OpenStudy (anonymous):

Evaluate the expression below when x = -4. 4x2 - 2x × 15

OpenStudy (anonymous):

A-56 b184 c376 D1,080

OpenStudy (anonymous):

anybody?

OpenStudy (anas.p):

\[4 \times 2 - 2(x) \times 15\] Is this the equation?

OpenStudy (anonymous):

4x2 - 2x × 15 and x= -4

OpenStudy (anonymous):

4x^2 - 2x times 15

OpenStudy (anas.p):

\[4x^2 - 2x \times 15\] this?

OpenStudy (anas.p):

Okay i think got it... You have to follow the BOD MAS rule. It will be easier if you solve it in this manner\[(4x^2) - ((2x) \times 15)\] Get it??.. now replace the x with (-4) If you still don't get it, ask.